# A Codimension Two Approach to The $\mathbb{S}^{1}$-Stability Conjecture

### v4

@@ 2. Technical Preliminaries
$\leqslant C_{4}(X,g)C_{5}(X,L,g)\lVert u\rVert_{\mathcal{L}^{2,p}(W)}\leqslant C_{1}(1+C_{1}^{\prime})C_{4}(X,g)C_{5}(X,L,g)\lVert F\rVert_{\mathcal{L}^{p}(W)}$
$<C_{1}(1+C_{1}^{\prime})C_{4}(X,g)C_{5}(X,L,g)\eta:=\eta^{\prime}.$
We use (11) in the next to last line of this equation. Shrinking $\eta$ as necessary finishes the proof. ∎
@@
(i) The only place we use the existence of a PSC metric on $X\times{\mathbb{S}}^{1}$ (i.e., $R_{h}>0$) is (8). The only place we use that $X$ and hence $W$ are oriented is doing the integration by parts/Stokes’ Theorem in (7).
@@
for $w\in W$ and fixed $P\in\mathbb{S}^{1}.$ Here $B_{1}(\phi_{W},g)=B_{1}(\phi_{W},g,P)$ is a globally defined first order operator on $u_{W}\in C^{\infty}(W)$; in particular, the left hand side of (17) has no derivatives in the $\mathbb{S}^{1}$-direction. Similarly, we have
@@
Take a chart $(x^{1},\ldots,x^{n},t)$ in $M$ containing a chart $(x^{1},\ldots,x^{n-1},t)$ in $W$ containing $P$, with $\theta=x^{n}$ the standard coordinate on $\mathbb{S}^{1}$ and $t:=x^{n+1}$ is the coordinate on $[0,1]$. Then
$\Delta_{g}u_{M}=g^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}-g^{ij}\Gamma_{ij,g}^{k}\frac{\partial u_{M}}{\partial x^{k}}.$
Here we sum over $i,j=1,\dotso,n+1$. Since $\partial u_{M}/\partial x^{n}=0$,
$g^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}=(\sigma^{*}g)^{ij}\frac{\partial^{2}u_{W}}{\partial x^{i}\partial x^{j}},$
where the sum is now over $i,j=1,\dotso,n-1,n+1$, so the second order terms in $\Delta_{g}u_{M}$ and $\Delta_{\sigma^{*}g}u_{W}$ agree. The Christoffel symbols $\Gamma^{k}_{ij,g}$ and $\Gamma^{k}_{ij,\sigma^{*}g}$ in the Laplacians may differ, but $\partial u_{M}/\partial x^{k}=\partial u_{W}/\partial x^{k}$. Thus the left hand side of (17) is a first order operator in $u_{W},$ so (17) follows.
@@
For fixed $\eta$ in Proposition 2.1, for fixed $P\in{\mathbb{S}}^{1}$ and $u_{W}$ as in Remark 2.1(ii), there exists a constant $K_{1}=K_{1}(\eta)>0$ such that
@@ 3. Proof of The Main Theorem
In this section, we prove the main theorem. One direction is trivial: if $X$ admits a PSC metric $g_{0}$, then the product metric on $X\times\mathbb{S}^{1}$ has PSC. We prove the other direction by our codimension two approach combined with conformal geometry.
As a possible approach to the ${\mathbb{S}}^{1}$-stability conjecture in the other direction, if we start with $R_{h}=R_{\imath^{*}g}>0$ on $X\times{\mathbb{S}}^{1}$, then $R_{\imath^{*}g}|_{X\times\{P\}}>0$. In the notation of (2), if we somehow knew that $R_{\imath^{*}g}|_{X\times\{P\}}-R_{\tau^{*}\imath^{*}g}|_{X}<0$ by some Gauss-Codazzi manipulations, then $R_{\tau^{*}\imath^{*}g}>0$ on $X$. Unfortunately, it is very hard to estimate the second order derivatives of the metric in the Gauss-Codazzi equation. Instead, we make two modifications: (i) we introduce the extra dimension $[-1,1]$ in order to produce the elliptic operator $L^{\prime}$ on $W$ in Lemma 2.1; (ii) motivated by [8], we consider a conformal transformation of $g$ to $\tilde{g}$. In terms of (2), the extra dimension in (i) allows us to transfer geometric information up the right hand side, i.e., from $(W,\sigma^{*}g)$ to $(M,g)$. The conformal transformation in (ii) produces $(M,\tilde{g})$ in the upper right corner, and Gauss-Codazzi moves us first to the upper left corner and then down to $(X,\tau^{*}\imath^{*}\tilde{g})$ in the lower left corner. In the end, we will prove that $\tau^{*}\imath^{*}\tilde{g}$ is a PSC metric on $X$, provided the $h$-angle between $\mu$ and $\partial_{\theta}$ is not too large on some $X\times\{P\}.$
@@
  • Pick $\eta\ll 1$ and $\delta=\delta(\eta)>0$ such that for the function $F=F(p,C,\delta)$ in Lemma 2.2, the solution $u_{W}$ of (14) satisfies $\frac{1}{2}<\|u_{W}\|_{\mathcal{C}^{0}}<\frac{3}{2}$; automatically, $\sup_{i,k}\|\partial_{x^{i}_{k}}u_{W}\|_{C^{0},\alpha}\leqslant\eta$ by Proposition 2.1 and the notation above the Proposition. We further shrink $\eta$ as necessary so that $\eta^{\prime}<1/4$ in (12).
  • Fix $P\in{\mathbb{S}}^{1}.$ Then $|B_{1}(u_{W},g,P)(w)|<K_{1}=K_{1}(\eta)$ by Lemma 2.5. Recall that $K_{1}\to 0$ as $\eta\to 0$, so shrinking the choice of $\eta$ above improves this estimate.
@@
  • Using the lower bound on $\|u_{W}\|$ above and $\|u_{W}\|_{\mathcal{C}^{1,\alpha}}<\eta$ by Proposition 2.1, we can choose $\eta\ll 1$ so that for all $x\in X,$
@@
We now work on the left side of (2) using the Gauss-Codazzi equation. Gauss-Codazzi and the conformal transformation law for the second fundamental form on $X\times\{P\}\subset X\times\mathbb{S}^{1}$ gives
@@
In the next equation, we use Remark 2.1(ii), Remark 2.1(iv), Lemma 2.3, and (18). In the first equality, the left hand side is evaluated at $x\in X$, and all terms on the right hand side (including $h_{\imath^{*}g}^{2}$ and $\lvert A_{\imath^{*}g}\rvert^{2}$) are evaluated at $y=(x,P)\in X\times\{P\}$; in the next inequality, the term $\Delta_{g}u_{M}$ is evaluated at $(x,0,P)\in M$; in the last equality, $K_{2}$ is evaluated at $x.$ We obtain
@@
The technical difficulty in the proof is assuring that the operator $L^{\prime}$ in Lemma 2.1 is elliptic, so that we get a smooth conformal factor $u_{W}.$ The difficult term $\nabla_{\mu}\nabla_{\mu}$ in $L^{\prime}$ appears because (i) we have to work on $M=X\times{\mathbb{S}}^{1}\times[0,1]$, as explained before Lemma 2.1, and (ii) to pass from $M$ to $X$, we have to use Gauss-Codazzi for $\tilde{g}$ twice:
@@
There are several known classes of counterexamples to the ${\mathbb{S}}^{1}$-stability conjecture, including odd degree hypersurfaces in $\mathbb{C}\mathbb{P}^{3}$ [7, Rmk. 1.25], and $M\ \#\ k\overline{\mathbb{C}\mathbb{P}^{2}}$ for a simply connected Kähler surface $M$ and a positive integer $k$ [5, Rmk. 5]. These are $4$-manifolds $X$ which admit no PSC metric, but such that $X\times S^{1}$ admits a PSC metric. For these and any other counterexamples, the PSC metric on $X\times S^{1}$ has large angles in the following sense:
@@
There is an analytic consequence of Theorem 3.1. Let $\lambda_{g}(X)$ be the Yamabe invariant of the closed manifold $X$ with respect to the conformal class $[g]$.

### v5

@@ 2. Technical Preliminaries
$\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g^{\prime})}$ $=\epsilon^{2-\frac{n}{2}}\left\lVert\frac{\partial^{2}u}{\partial t^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g)}\leqslant\epsilon^{2-\frac{n}{2}}\left\lVert u\right\rVert_{H^{2}(V_{2},\sigma^{*}g)}$
(19) $\leqslant\epsilon^{2-\frac{n}{2}}D_{1}\left(\lVert L_{g}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}+\left\lVert u\right\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}\right)$
$\leqslant\epsilon^{2-\frac{n}{2}}D_{1}\left(\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}+D_{1}^{\prime}\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g)}\right)$
$\leqslant D_{1}\epsilon^{2}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime}))}+D_{1}D_{1}^{\prime}\epsilon^{2-\frac{n}{n+1}}\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g^{\prime})}.$
Since $g=h\oplus dt^{2}$, the mean curvature satisfies $h_{g}\equiv 0$. Also, since $R_{g}>0$ and $\dim M=n+1$, the relative Yamabe constant [2] $\lambda:=\lambda(M,\partial M,[g])$ for the conformal class $[g]$ on $(M,\partial M)$ is positive and satisfies
(20) $\lVert v\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(M,g)}^{2}\leqslant\lambda^{-1}\left(\frac{4n}{n-1}\lVert\nabla_{g}v\rVert_{\mathcal{L}^{2}(M,g)}^{2}+\int_{M}R_{g}v^{2}{\rm dvol}_{g}\right),\forall v\in H^{1}(M,g).$
$\lambda$ is a conformal invariant, so we can replace $g$ by $g^{\prime}$ in (20).
Let $\bar{u}=u\circ\pi:M\to\mathbb{R}$ be the pullback of $u:W\to\mathbb{R}$ under the projection $\pi:M\to W$, so $\bar{u}$ is constant on the $t$-fibers. Then for $\theta^{\prime}=\epsilon^{-1}\theta,$
[…]
where $D_{2}$ satisfies $\text{dvol}_{d(\theta^{\prime})^{2}\oplus\sigma^{*}g^{\prime}}\leq D_{2}^{\frac{2(n+1)}{n-1}}\ \text{dvol}_{g^{\prime}}$. The fact that $D_{2}$ is independent of $\epsilon$ (and $u$) is the reason we scale $\theta.$
Using (20) for $g^{\prime}$, (21) becomes
${\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g^{\prime})}}$
$\leqslant\epsilon^{\frac{n-1}{2(n+1)}}D_{2}\lambda^{-1/2}\left(\frac{4n}{n-1}\lVert\nabla_{g^{\prime}}\bar{u}\rVert^{2}_{\mathcal{L}^{2}(M,g^{\prime})}+\int_{M}R_{g^{\prime}}\bar{u}^{2}\text{dvol}_{g^{\prime}}\right)^{\frac{1}{2}}$
(22) $\leqslant\epsilon^{\frac{n-1}{2(n+1)}}D_{2}D_{3}\lambda^{-1/2}\left(\frac{4n}{n-1}\cdot\frac{1}{2\pi}\int_{\mathbb{S}^{1}}d\theta^{\prime}\int_{W}\lvert\nabla_{\sigma^{*}g^{\prime}}u\rvert^{2}_{\sigma^{*}g^{\prime}}\text{dvol}_{\sigma^{*}g^{\prime}}\right.$
$\qquad\left.+\frac{1}{2\pi}\int_{\mathbb{S}^{1}}d\theta^{\prime}\int_{W}R_{g^{\prime}}u^{2}\text{dvol}_{\sigma^{*}g^{\prime}}\right)^{\frac{1}{2}}$
$=D_{2}D_{3}\lambda^{-1/2}\left(\frac{4n}{n-1}\lVert\nabla_{\sigma^{*}g^{\prime}}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}^{2}+\epsilon^{2}\max_{W}\lvert R_{g}\rvert\lVert u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}\right)^{\frac{1}{2}},$
where we have $\text{dvol}_{g^{\prime}}\leq D_{3}^{2}\text{dvol}_{d(\theta^{\prime})^{2}\oplus\sigma^{*}g^{\prime}}$. We emphasize that the introduction of $\bar{u}$ lets us use the positivity of $R_{g}$ and hence $\lambda$ on $M$ to get the estimate (2) for any function $u$ on $W$.
We also have
(23) $\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{C}^{0}(V_{1})}\leqslant D_{0}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{H^{s}(V_{1},\sigma^{*}g)}=D_{0}\epsilon^{\frac{n}{2}}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{H^{s}(V_{1},\sigma^{*}g^{\prime})}.$
The first inequality uses the Sobolev embedding theorem. For the second inequality, the zeroth order contribution to the $H^{s}$ norm contributes $\epsilon^{\frac{n}{2}}$ from the scaling of the volume form; higher order derivatives in $t$ contribute higher powers of $\epsilon,$ which can be ignored.
Applying (8), (18), (2), (2) to (23), we get
${\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{C}^{0}(V_{1})}\leqslant D_{0}D_{s}\epsilon^{\frac{n}{2}}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g^{\prime})}}$
$\leqslant D_{0}D_{1}D_{1}^{\prime}D_{s}\epsilon^{2-\frac{n}{n+1}+\frac{n}{2}}\lVert u\rVert_{\mathcal{L}^{\frac{2(n+1)}{n-1}}(W,\sigma^{*}g^{\prime})}+D_{0}D_{1}D_{s}\epsilon^{2+\frac{n}{2}}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}$
$\leqslant D_{0}D_{1}D_{1}^{\prime}D_{2}D_{3}D_{s}\lambda^{-1/2}\epsilon^{2-\frac{n}{n+1}+\frac{n}{2}}\left(\frac{4n}{n-1}\lVert\nabla_{\sigma^{*}g^{\prime}}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}^{2}+\epsilon^{2}\max_{W}\lvert R_{g}\rvert\lVert u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}\right)^{\frac{1}{2}}$
$\qquad+D_{0}D_{1}D_{s}\epsilon^{2+\frac{n}{2}}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g^{\prime})}$
$=D_{0}D_{1}D_{1}^{\prime}D_{2}D_{3}D_{s}\lambda^{-1/2}\epsilon^{2-\frac{n}{n+1}+1}\left(\frac{4n}{n-1}\lVert\nabla_{\sigma^{*}g}u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}^{2}+\max_{W}\lvert R_{g}\rvert\lVert u\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}\right)^{\frac{1}{2}}$
$\qquad+\epsilon^{2}\lVert F\rVert_{\mathcal{L}^{2}(W,\sigma^{*}g)}$
$<\bar{D}\epsilon^{2+\frac{1}{n+1}}\eta+\bar{D}^{\prime}\epsilon^{2}(C+1)\epsilon^{1/2},$
for positive constants $\bar{D},\bar{D}^{\prime}$. In the last line, $\eta$ comes from Prop. 2.1, with $C$ fixed and $\epsilon$ in this proposition to be determined; for the last term, we use $\text{supp}(F)\subset X\times[-\epsilon,\epsilon]$ in Lem. 2.2. If the estimate in this lemma holds for a fixed $C$ and $\epsilon_{0}$, then it holds for all $\epsilon<\epsilon_{0}.$ Therefore,
$\left\lVert\frac{\partial^{2}u}{\partial t^{2}}\right\rVert_{\mathcal{C}^{0}(U_{1})}=\epsilon^{-2}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{C}^{0}(V_{1})}<\bar{D}\epsilon^{\frac{1}{n+1}}\eta+\bar{D}^{\prime}(C+1)\epsilon^{1/2}<\eta^{\prime},$
by fixing $\eta$ and shrinking $\epsilon$ as needed. ∎
@@
(i) The only places in the paper we use the existence of a PSC metric $h$ on $X\times{\mathbb{S}}^{1}$ are in (10) and in the positivity of the relative Yamabe invariant above (20). The only place we use that $X$ and hence $W$ are oriented is doing the integration by parts/Stokes’ Theorem in (2).
@@
for $w\in W$ and fixed $P\in\mathbb{S}^{1}.$ Here $B_{1}(u_{W},g)=B_{1}(u_{W},g,P)$ is a globally defined first order operator on $u_{W}\in C^{\infty}(W)$; in particular, the left hand side of (28) has no derivatives in the $\mathbb{S}^{1}$-direction. Similarly, we have
@@
In (28), for $\pi:M\to W$, we are computing
(30) $\Delta_{g}u_{M}-\pi^{*}(\Delta_{\sigma^{*}g}u_{W})=\Delta_{g}(u_{W}\circ\pi)-\pi^{*}(\Delta_{\sigma^{*}g}(u_{W}))=\Delta_{g}(u_{W}\circ\pi)-\pi^{*}(\Delta_{\sigma^{*}g}(u_{M}\circ\sigma))$
in a tubular neighborhood $T$ of the hypersurface $H:=W\times\{P\}\subset M$, and then restricting to $H$. By taking the level surfaces for the $g$-distance function to $H$, we can take a chart $(x^{1},\ldots,x^{n},t)$ in $M$ containing a chart $(x^{1},\ldots,x^{n-1},t)$ for $W\times\{P\}$, with $t:=x^{n+1}$ the coordinate on $[-1,1]$ and $\partial_{x^{n}}$ normal to the level surfaces. If $x^{n}(P)=0$ in these coordinates, we can write $h=h_{X,x_{n}}\oplus d(x^{n})^{2}$ and $g=h_{X,x_{n}}\oplus d(x^{n})^{2}\oplus dt^{2}$ near $W\times\{P\}$. Therefore, $\sigma^{*}g=h_{X,0}\oplus dt^{2}$ on $W\approx W\times\{P\}$, and $g^{ij}=(\sigma^{*}g)^{ij},i,j=1,\dotso,n-1$ on $W$. In summary, on $T$,
$(g_{ij})=\begin{pmatrix}g_{11}&\dotso&g_{1,n-1}&0&0\\ \vdots&\dotso&\vdots&\vdots&\vdots\\ g_{n-1,1}&\dotso&g_{n-1,n-1}&0&0\\ 0&\dotso&0&1&0\\ 0&\dotso&0&0&1\end{pmatrix},$
with $(g^{ij})$ the inverse of the whole matrix, and $(\sigma^{*}g)^{ij}$ the inverse of the minor given by deleting the $n^{\rm th}$ row and column. Then
$\Delta_{g}u_{M}=\sum_{i,j=1}^{n-1}g^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}+\frac{\partial^{2}u_{M}}{\partial(x^{n})^{2}}+\frac{\partial^{2}u_{M}}{\partial t^{2}}-\sum_{i,j,k=1}^{n+1}g^{ij}\Gamma_{ij,g}^{k}\frac{\partial u_{M}}{\partial x^{k}}.$
The projection map $\pi:M\rightarrow W$ is $\pi(x^{1},\dotso,x^{n-1},\theta,t)=(x^{1},\dotso,x^{n-1},t)$, so $\frac{\partial u_{M}}{\partial x^{n}}=\frac{\partial(u_{W}\circ\pi)}{\partial x^{n}}=0.$ Similarly, it is easy to check that
$\frac{\partial u_{W}}{\partial x^{i}}\biggl{|}_{w}=\frac{\partial(u_{M}\circ\sigma)}{\partial x^{i}}\biggl{|}_{w}=\frac{\partial u_{M}}{\partial x^{i}}\biggl{|}_{(w,P)},i=1,\dotso,n-1,n+1.$
On $W$,
$\Delta_{\sigma^{*}g}u_{W}=\sum_{i,j=1}^{n-1}(\sigma^{*}g)^{ij}\frac{\partial^{2}u_{W}}{\partial x^{i}\partial x^{j}}+\frac{\partial^{2}u_{W}}{\partial t^{2}}-(\sigma^{*}g)^{ij}\Gamma_{ij,\sigma^{*}g}^{k}\frac{\partial u_{W}}{\partial x^{k}}.$
Since $(g^{ij})=(\sigma^{*}g)^{ij}$ on $W$, we have
$\Delta_{g}u_{M}(w,P)-\Delta_{\sigma^{*}g}u_{W}(w)$
$\qquad=\left(\sum_{i,j=1}^{n-1}g^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}\biggl{|}_{(w,P)}+\frac{\partial^{2}u_{M}}{\partial t^{2}}\biggl{|}_{(w,P)}-\sum_{i,j,k=1}^{n+1}g^{ij}\Gamma_{ij,g}^{k}\frac{\partial u_{W}}{\partial x^{k}}\biggl{|}_{w}\right)$
$\qquad\qquad-\left(\sum_{i,j=1}^{n-1}(\sigma^{*}g)^{ij}\frac{\partial^{2}u_{M}}{\partial x^{i}\partial x^{j}}\biggl{|}_{(w,P)}+\frac{\partial^{2}u_{M}}{\partial t^{2}}\biggl{|}_{(w,P)}-\sum_{i,j,k\neq n}(\sigma^{*}g)^{ij}\Gamma_{ij,\sigma^{*}g}^{k}\frac{\partial u_{W}}{\partial x^{k}}\biggl{|}_{w}\right)$
$\qquad:=B_{1}(u_{W},g)(w).$
@@
For fixed $\eta$ in Prop. 2.1, for fixed $P\in{\mathbb{S}}^{1}$ and $u_{W}$ as in Remark 2.1(ii), there exists a constant $K_{1}=K_{1}(\eta)>0$ such that
@@ 3. Proof of The Main Theorem
In this section, we prove Thm.  3.1. One direction of the ${\mathbb{S}}^{1}$-Stability Conjecture direction is trivial: if $X$ admits a PSC metric $g_{0}$, then the product metric on $X\times\mathbb{S}^{1}$ has PSC.
As a possible approach to the ${\mathbb{S}}^{1}$-stability conjecture in the other direction, if we start with $R_{h}=R_{\imath^{*}g}>0$ on $X\times{\mathbb{S}}^{1}$, then $R_{\imath^{*}g}|_{X\times\{P\}}>0$. In the notation of (2), if we somehow knew that $R_{\imath^{*}g}|_{X\times\{P\}}-R_{\tau^{*}\imath^{*}g}|_{X}<0$ by some Gauss-Codazzi manipulations, then $R_{\tau^{*}\imath^{*}g}>0$ on $X$. Unfortunately, it is very hard to estimate the second order derivatives of the metric in the Gauss-Codazzi equation. Instead, we make two modifications: (i) we introduce the extra dimension $[-1,1]$ in order to produce the elliptic operator $L^{\prime}$ on $W$ in Lem. 2.1; (ii) motivated by [9], we consider a conformal transformation of $g$ to $\tilde{g}$ as in (25). In terms of the diagram (2), the extra dimension in (i) allows us to transfer geometric information up the right hand side, i.e., from $(W,\sigma^{*}g)$ to $(M,g)$. The conformal transformation in (ii) produces $(M,\tilde{g})$ in the upper right corner, and Gauss-Codazzi moves us first to the upper left corner and then down to $(X,\tau^{*}\imath^{*}\tilde{g})$ in the lower left corner. In the end, we will prove that $\tau^{*}\imath^{*}\tilde{g}$ is a PSC metric on $X$, provided (1) holds.
@@
  • Pick $\eta\ll 1$ and $\delta=\delta(\eta)>0$ such that for the function $F=F(p,C,\delta)$ in Lem. 2.2, the solution $u_{W}$ of (24) satisfies $\frac{1}{2}<\|u_{W}\|_{\mathcal{C}^{0}}<\frac{3}{2}$; automatically, $\sup_{i,k}\|\partial_{x^{i}_{k}}u_{W}\|_{C^{0,\alpha}}\leqslant\eta$ by Prop. 2.1 and the notation above the Proposition. We further shrink $\eta$ as necessary so that $\eta^{\prime}<1/4$ in (14).
  • Fix $P\in{\mathbb{S}}^{1}.$ Then $|B_{1}(u_{W},g,P)(w)|<K_{1}=K_{1}(\eta)$ by Lem. 2.5. Recall that $K_{1}\to 0$ as $\eta\to 0$, so shrinking the choice of $\eta$ above improves this estimate.
@@
    Here $h_{\imath^{*}g},A_{\imath^{*}g}$ are the mean curvature and second fundamental form of $i^{*}g.$
  • Using the lower bound on $\|u_{W}\|$ above and $\|u_{W}\|_{\mathcal{C}^{1,\alpha}}<\eta$ by Prop. 2.1, we can choose $\eta\ll 1$ so that for all $x\in X,$
@@
We now work on the left side of (2) using the Gauss-Codazzi equation. Gauss-Codazzi and the conformal transformation law for the second fundamental form on $X\times\{P\}\subset X\times\mathbb{S}^{1}$ give
@@
In the next equation, we use Remark 2.1(ii), Remark 2.1(iv), Lem. 2.3, and (29). In the first equality, the left hand side is evaluated at $x\in X$, and all terms on the right hand side (including $h_{\imath^{*}g}^{2}$ and $\lvert A_{\imath^{*}g}\rvert^{2}$) are evaluated at $y=(x,P)\in X\times\{P\}$; in the next inequality, the term $\Delta_{g}u_{M}$ is evaluated at $(x,0,P)\in M$; in the last equality, $K_{2}$ is evaluated at $x.$ We obtain
@@
The technical difficulty in the proof is assuring that the operator $L^{\prime}$ in Lem. 2.1 is elliptic, so that we get a smooth conformal factor $u_{W}.$ The difficult term $\nabla_{\mu}\nabla_{\mu}$ in $L^{\prime}$ appears because (i) we have to work on $M=X\times{\mathbb{S}}^{1}\times[0,1]$, as explained before Lem. 2.1, and (ii) to pass from $M$ to $X$, we have to use Gauss-Codazzi for $\tilde{g}$ twice:
@@
There are several known classes of counterexamples to the ${\mathbb{S}}^{1}$-stability conjecture, including odd degree hypersurfaces in $\mathbb{C}\mathbb{P}^{3}$ [8, Rmk. 1.25], and $M\ \#\ k\overline{\mathbb{C}\mathbb{P}^{2}}$ for a simply connected Kähler surface $M$ and a positive integer $k$ [6, Rmk. 5]. These are $4$-manifolds $X$ which admit no PSC metric, but such that $X\times S^{1}$ admits a PSC metric. For these and any other counterexamples, the PSC metric on $X\times S^{1}$ has large angles in the following sense:
@@
There is an analytic consequence of Thm.  3.1. Let $\lambda_{g}(X)$ be the Yamabe invariant of the closed manifold $X$ with respect to the conformal class $[g]$.